OpenStudy (kaylaprincess):

help as soon as possible : what would the standard form be for this equation? f(x) = 3(x-1)^2+24

OpenStudy (zale101):

The standard form would be: f(x)=ax^2+bx+c To make your parabola similar to that standard equation, you then need to simplify f(x) = 3(x-1)^2+24.

OpenStudy (kaylaprincess):

f(x) = 3(x-1)(x-1)^2+24 is that how its supposed to start? ( iwasnt too sure )

OpenStudy (campbell_st):

not quite inside the brackets is a perfect square so it becomes \[f(x) = 3(x^2 -2x +1) + 24\] just distribute and collect like terms

OpenStudy (kaylaprincess):

f(x) = 3x^2 - 6x + 3 + 24

OpenStudy (kaylaprincess):

f(x) = 3x^2 -6x +27

OpenStudy (campbell_st):

that's correct

OpenStudy (kaylaprincess):

could you help me with one more? Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. Basically you just plug x into the function and solve from there right? ( ill give you another medal on someone elses post )

OpenStudy (campbell_st):

well you need to plug them in and find f(4) and f(7) then jus find the slope between the ordered pairs (4, f(4)) and (7, f(7)) hope it helps

OpenStudy (kaylaprincess):

into the equation ?